OpenStudy (anonymous):

evaluate the discriminant of this equation and determine how many real solutions it has x^2+5x+6=0

OpenStudy (anonymous):

Use factorization to find the roots.

OpenStudy (anonymous):

(x+3)(x+2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I factored the equation for you which is most of the work, but in order to find the factors you need to find two numbers when multiplied that equal 6 and two numbers when added or subtracted that equal 5.

OpenStudy (anonymous):

the numbers must be the same for both relationships, and those two numbers will be part of the factorized solution as I've written above.

OpenStudy (anonymous):

please can u just give me the answer i dont know

OpenStudy (anonymous):

I gave you most of the answer, you should try and look at your notes to finish the problem.

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

HCfootball do you know the quadratic formula? it lokk lik this \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] if you do use it to get the answers, if not disregard this reply

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

but i still can figure it out

OpenStudy (anonymous):

ok so x^2+5x+6=0 a is always the number that multiplies by x^2 in this case it's 1 b is always the number that multiplies by x, it's 5 c is the the on that dosn't have an x can you try it now , it will give you 2 answers

OpenStudy (anonymous):

\[b ^{2} - 4ac\] is the discriminant if it's less than zero, the quadratic has no real roots if it's equal to zero, the quadratic has one real root if it's greater than zero, the quadratic has two real roots

OpenStudy (anonymous):

final awnser x=-3 and x=-2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

(x+3)(x+2) (x+3)=0 x+3=0 x=-3 (x+2)=0 x+2=0 x=-2